OpenStudy (rayman5766):

What is the value of the expression when x = –26?

OpenStudy (rayman5766):

What is the value of the expression x -6 /4 when x = –26?

OpenStudy (rayman5766):

What is the value of the expression x -6 /4 when x = –26?

geerky42 (geerky42):

\(\dfrac{x-6}{4}\), right? Just replace x to -26 and solve.